you are doing a statistical experiment where you roll a die and pick a color. the color choices are blue, green, or purple. Give the sample space for this experiment. What is the probability that the outcome of the experiment is:
a) a purple and a 3?
b) a blue and an even number?
c) a green or a 5?

I greatly appreciate all your help with this. thank you!

Answer by ewatrrr(24785)


Hi
statistical experiment where you roll a die and pick a color
die has six possible outcomes(3 of which are even)
color 3 possible outcomes
P(a purple 'and' a 3) = 1/3*1/6 = 1/18
P(a blue 'and' an even number) = 1/3*3/6 = 3/18 = 1/6
P(a green 'or' a 5) = 1/3 + 1/6 - 1/18 = 2/6 + 1/6 - 1/18 = 9/18 - 1/18 = 8/18